Circular No. 6986

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1998 P1 D. A. J. Seargent, The Entrance, New South Wales, Australia, reports

the visual discovery by Peter Williams of Heathcote (near Sydney) of a new comet. Available observations:

1998 UT  α_{2000}  δ_{2000}  m_{1} Observer
Aug. 10.486  15^{h}10.^{m}8  −64^{o}59′  9.5 Williams
11.37962 15 02 11.23 −63 54 16.7  Garradd
11.38046 15 02 10.74 −63 54 13.4  _{”}
11.38200 15 02 09.89 −63 54 06.7  _{”}
11.38350 15 02 08.99 −63 53 59.3  _{”}
11.385 15 02.3 −63 54 9.5 Williams
11.38775 15 02 06.78 −63 53 42.1  Garradd
11.46470 15 01 25.34 −63 48 12.9  _{”}
11.46569 15 01 24.81 −63 48 08.5  _{”}
11.47476 15 01 19.94 −63 47 29.6  _{”}
11.48030 15 01 17.02 −63 47 05.7  _{”}
11.48244 15 01 15.86 −63 46 56.5  _{”}

P. Williams (Heathcote). 0.30-m f/6 reflector (72×). Comet diffuse with a

stellar central condensation of mag 13. G. J. Garradd (Loomberah). 0.45-m f/5.4 reflector + CCD. 60-sec exposure

shows 8′ coma and 6′ tail in p.a. 103^{o}; m_{2} = 13.2–14.4.

Additional visual m_{1} estimates: Aug. 10.52 UT, 9.5 (Seargent, 25×100 binoculars; poor conditions, moonlight); 11.38, 9.5 (Williams, 20×80 binoculars; coma diameter 4′); 11.44, 9.0: (R. H. McNaught, Siding Spring Observatory, 20×120 binoculars).

SUPERNOVA 1998dj in NGC 788 M. Modjaz, T. Shefler, E. Halderson, J. Y. King, W. D. Li, R. R.

Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent supernova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1998dj was found and confirmed on unfiltered CCD images obtained on Aug. 8.5 (mag ∼ 16.1) and 9.5 UT (mag ∼ 16.0). The object is located at α = 2^{h}01^{m}06^{s}.92, δ = −6^{o}49′04′′.8 (equinox 2000.0), which is ∼ 7′′.1 east and 8′′.6 south of the nucleus of NGC 788. A KAIT image (limiting mag ∼ 18.0) obtained on Jan. 25.2 does not show a star at the position of the new object.

1998 August 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green